OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. y - 2 = 4(x + 6) y + 6 = 4(x - 2) y + 2 = 4(x - 6) y - 6 = 4(x + 2)

OpenStudy (anonymous):

Point slope form is: y-y1=m(x-x1)

OpenStudy (anonymous):

You just have to plug in the point and the slope :)

OpenStudy (anonymous):

Use:\[y-y1=m(x-x1)\]Substituting (x1=-2, y1=6, m=4)\[y-6=4(x-(-2))\]So the answer is:\[y-6=4(x+2)\]